OpenStudy (anonymous):

-4 (x+10)-6=-3 (x-2)

OpenStudy (anonymous):

Your goal is to get x on one side and the numbers on the other. PEMDAS is always important in questions like this. On the left side, we see that -4 can be multiplied by (x+10), which we're going to do because our goal is to simplify each side - it makes it easier. So multiply -4 by (x+10), leaving you with: -4x-40-6=-3(x-2) You can continue to simplify the left side: -4x-46=-3(x-2) Then, when you can simplify the left no longer, move to the right: -4x-46=-3x+6 Now, like I said, move the x to the left of the equation, and the constants (the numbers without a variable) to the right. -x=52 Now, when you get a negative sign on the x side, you need to get rid of it, because you want the value of positive x. To do this, just flip the signs on each side of the equation. You'll always be allowed to change the signs of an equation as long as you do it on both sides. So... x=-52 That's it! Any questions?

OpenStudy (anonymous):

x=-52